I have worked on this problem and I do not know where I'm going wrong but I know the answer I'm coming up with is incorrect.
Use elimination to solve: x/3+4y=30 and 2x-y=5 I got 22/25 and -2 I just know this is incorrect. PLEASE HELP!!!!
here is another one I can not figure out: Use elimination to solve: 6x-y/4=-9
and 3x+2y= -30
this would be a great help. thank you :)

Answer by daleswift(2)

x/3 + 4y = 30
2x - y = 5
Solve for x: To eliminate the y's, multiply each side of second equation by 4 (you could also multiply each side of the first equation by 1/4)
-----------
so, the second equation becomes:
4 [2x - y] = 4 [5]
or
8x - 4y = 20
now, eliminate the y's and simplify:
x/3 + 8x = 50
3 [x/3 + 8x] = 3 [50]
x + 24x = 150
25x = 150
x = 6
------------------------------------
Now, solve for y by eliminating the x's. Do this by multiplying each side of the first equation by "-6" (or, you could also multiply both sides of the second equation by "-1/6")
so,
if you multiply both sides of the first equation by "-6", you will have:
(-6) [x/3 + 4y] = (-6) [30]
or
-2x - 24y = -180
Now, the two equations can be stated as:
1) -2x - 24y = -180
2) 2x - y = 5
Now, combine the equations to eliminate the x's:
-25y = -175 (if you now divide each side by "-25" you have):
y = 7
Therefore, your answer is (6,7) you can check your answer by plugging (6,7) into the two original equations.